# Geodesic Distance Between Graphs: A Spectral Metric for Assessing the Stability of Graph Neural Networks

### v1

@@ A.4 Experimental Setup
To evaluate the performance of the Graph Geodesic Distance (GGD) metric, we utilized graph datasets from the TUDataset collection Morris et al. [2020]. For small graphs, we used datasets like MUTAG and BZR, and for larger graphs, we selected PC-3H and SW-620H, which present more sizable networks.
@@
While Classification tasks, each dataset was split into 90% training and 10% testing sets to ensure an unbiased evaluation process. When assessing the correlation with GNN, we trained a 3-layer GIN with 90% of all graphs from MUTAG and validated with the rest 10%. For the performance evaluation using graphs with partial node features, we took each dataset with node features and randomly removed a certain portion of features.
@@
All experiments have been evaluated on a laptop with an Apple M1 chip, featuring an 8-core CPU and a 7-core GPU.

### v2

@@ A.10 Experimental Setup
To evaluate the performance of the Graph Geodesic Distance (GGD) metric, we utilized graph datasets from the TUDataset collection (Morris et al., 2020). For small graphs, we used datasets like MUTAG and BZR, and for larger graphs, we selected PC-3H and SW-620H, which present more sizable networks. Detailed information about the datasets used is provided in Table 10.
Table 10: Brief description of graph datasets used.
Dataset name Number of graphs Average number of nodes Average number of edges
MUTAG 188 17.93 19.79
PC-3H 27509 47.20 49.33
SW-620H 40532 26.06 28.09
BZR 405 35.75 38.36
@@
While Classification tasks, each dataset was split into 90% training and 10% testing sets to ensure an unbiased evaluation process. When assessing the correlation with GNN, we trained a three-layer GIN with 90% of all graphs from MUTAG and validated with the rest 10%. For the performance evaluation using graphs with partial node features, we took each dataset with node features and randomly removed a certain portion of features.
@@
All experiments have been evaluated on a laptop with an Apple M1 chipset, featuring an eight-core CPU and a seven-core GPU.